# Coresets from Trajectories: Selecting Data via Correlation of Loss Differences

### v1

@@ 7 Computational and Storage Efficiency
Beyond accuracy, a practical coreset method should keep both compute and storage costs low. CLD does so by relying only on per-sample loss scalars that standard training already produces; no per-sample gradients, Hessians, or pairwise similarities are required. The only extra work is a forward-only sweep over a small held-out query set each proxy epoch ($Q\ll N$) to record query losses. In contrast, gradient/adversarial methods incur extra backward passes, while similarity/nearest-neighbor methods require feature extraction and large feature caches. We quantify the end-to-end compute cost (selection plus training on the selected coreset) and the storage overhead in detail in Appendix˜E and summarize the symbolic complexity below.
@@ 8 Discussion
Several methods incorporate bias-reduction mechanisms, such as $\mathtt{CCS}$ (Zheng et al., 2022), which stratifies selection across score percentiles to promote diversity. In contrast, $\mathtt{CLD}$ leverages per-class validation trajectories, yielding a generalization-aware signal that naturally balances classes and downweights noisy or redundant samples. As shown in Figure˜5(b), applying $\mathtt{CCS}$-style stratified sampling on top of $\mathtt{CLD}$ scores consistently reduces accuracy across coreset sizes on CIFAR-100. This contrasts with metrics like $\mathtt{AUM}$, where $\mathtt{CCS}$ can be beneficial; for $\mathtt{CLD}$, percentile quotas perturb its validation-aligned ranking and reintroduce less informative points.
@@
Our theoretical guarantees for $\mathtt{CLD}$ require that the validation set be a faithful proxy for the test distribution (Assumption˜3 in Section˜5). To understand this, we ask: How sensitive is $\mathtt{CLD}$ to the composition of the validation set, and does bias in this validation set affect downstream coreset quality? To probe this, we split CIFAR-100’s $50\text{k}$ training set into a classwise $25\text{k}/25\text{k}$ train/pool partition. From the pool, we constructed $5000$ example validation sets using five heuristics based on publicly available memorization scores (Feldman & Zhang, 2020), denoted $\mathtt{mem}$. Intuitively, low-$\mathtt{mem}$ points correspond to canonical, stereotypical examples (often helpful for transfer and exploited by methods such as SloCurv), while high-$\mathtt{mem}$ points surface atypical or mislabeled examples. These heuristics, therefore, let us bias the validation set toward “typical” or “atypical” regions of the data. We trained ResNet-18 proxies on the $25\text{k}$ train split, computed $\mathtt{CLD}$ with respect to each heuristic-based validation set, built class-balanced coresets, and fine-tuned ResNet-18 across coreset sizes (five seeds; see Figure˜6(b)). As shown in Figure˜6(a), the non-random heuristics yield validation sets with markedly different $\mathtt{mem}$ profiles. Two clear patterns emerge. First, validation sets biased toward Highest-$\mathtt{mem}$ examples consistently degrade performance across coreset sizes: overrepresenting atypical or mislabeled samples lowers accuracy and hinders learning. Second, validation sets built from Lowest-$\mathtt{mem}$ examples generally support stronger generalization, but retaining some high-$\mathtt{mem}$ points is beneficial for capturing long-tail behavior likely present in the test distribution. In our runs, Proportional sampling, drawing examples according to the pool’s original $\mathtt{mem}$ distribution, was the most reliable overall, and would likely improve further if mislabeled points were filtered or downweighted (which we did not do in this study). These findings underscore that $\mathtt{CLD}$’s effectiveness depends critically on the quality of the validation signal: coresets cannot exceed the fidelity of the validation dynamics they are aligned with. This motivates future work on principled procedures to build clean, representative validation sets (e.g., robust to label noise) and to select validation sizes that balance reliability with efficiency.
@@ Acknowledgements
This work was supported in part by the Center for the Co-Design of Cognitive Systems (CoCoSys), a DARPA-sponsored JUMP 2.0 center, the Semiconductor Research Corporation (SRC), the National Science Foundation, and Collins Aerospace. We are also thankful to Efstathia Soufleri, Amitangshu Mukherjee, Utkarsh Saxena, and Sakshi Choudhary for their helpful discussions and feedback.
@@ Appendix B Detailed Theoretical Framework
B.1 Supporting Lemmas for Theorem˜1
@@
B.2 Proof for Theorem˜1
@@ Appendix C Datasets, Models, and Experimental Details
We evaluate our method on two standard image classification benchmarks: CIFAR-100 (Krizhevsky et al., 2009) and ImageNet-1k (Russakovsky et al., 2015). CIFAR-100 consists of 50,000 training and 10,000 test images across 100 classes and is publicly available without licensing restrictions. For ImageNet-1k, we use the official release from https://image-net.org/download.php, which is provided under a standard academic research license and requires user agreement to the terms of access.
@@
For baseline comparisons, we use the DeepCore library Guo et al. (2022) (https://github.com/PatrickZH/DeepCore), which provides standardized implementations of several coreset selection techniques and is licensed under MIT. All training and evaluation code was implemented in PyTorch; dependencies including torchvision are MIT/BSD licensed.
@@
All networks were trained using SGD Bottou (2010) for $164$ epochs with an initial learning rate of $0.1$, decayed by a factor of $0.1$ at epochs $81$ and $121$. Nesterov momentum Sutskever et al. (2013) with momentum $0.9$ was used, along with weight decay $5\times 10^{-4}$. Standard augmentations included resizing to $32\times 32$, random cropping with padding =4=4, random horizontal flips, and normalization.
@@ Appendix E Detailed Explanation of Compute and Storage Cost of Coreset Methodologies
  • We use the standard ImageNet recipe of $T{=}T_{\text{proxy}}{=}90$ epochs (Yang, 2017).
@@ E.1.1 Kernel Herding (Herding)
Herding (Chen et al., 2010) iteratively constructs a representative subset by approximating the data distribution in an RKHS. At iteration $t$, it selects
@@ E.1.2 Example Forgetting (Forgetting)
Forgetting (Toneva et al., 2018) measures, for each training sample, how many times it transitions from being correctly classified to incorrectly classified during training (“forgetting events”). Examples with higher forgetting counts are ranked as more informative.
@@ E.1.3 Area Under Margin (AUM)
AUM (Pleiss et al., 2020) scores each training sample by aggregating its margin over training (e.g., logit of the true class minus the max non-true logit), producing the area under the margin across epochs/updates. Higher absolute AUM indicates more consistently confident predictions; lower AUM can flag ambiguous or noisy samples. We compute AUM with a proxy model and then train the large model on the selected coreset.
@@ E.1.4 Contrastive Active Learning (Cal)
Cal (Margatina et al., 2021) acquires unlabeled examples that are near labeled ones in feature space yet differ in predictive probabilities (contrastive pairs), using nearest neighbors over encoder features and a simple divergence-based ranking.
@@ E.1.5 Gradient and Error L2 Norm-based Data Pruning (GraNd, EL2N)
GraNd (Paul et al., 2021) ranks training examples by the (expected) per-example gradient norm early in training:
@@
EL2N (Paul et al., 2021) ranks examples by the (expected) L2 error of predictions early in training:
@@ E.1.6 Using Class Feature Medians (Moderate)
Moderate (Xia et al., 2022) builds a representative coreset by selecting, within each class, the samples whose feature-to-center distances are closest to that class’s median distance (thus avoiding both easy near-center redundancies and far-out outliers). We compute class centers and distances in a proxy feature space.
@@ E.1.7 Message Passing ($\mathbb{D}^{2}-\mathtt{Pruning}$)
$\mathbb{D}^{2}\mathtt{Pruning}$ (Maharana et al., 2024) selects a coreset by balancing difficulty and diversity via message passing on a dataset graph built from proxy features. Initial per-sample difficulty scores (from the proxy) are diffused over a kkNN graph so that each example’s score incorporates information from its neighbors; a graph-based sampler then selects a subset that covers diverse yet difficult regions.
@@ E.2.1 Gradient Matching Optimization (CRAIG)
CRAIG (Mirzasoleiman et al., 2020) selects a subset whose (aggregated) gradients closely match those of the full dataset across training, typically using a submodular (stochastic-greedy) objective over per-sample gradient embeddings at selected anchor epochs. We compute embeddings with a proxy model and then train the large model on the coreset for all $T$ epochs.
@@ E.2.2 Generalization-based Data Subset Selection for Efficient and Robust Learning (Glister)
Glister (Killamsetty et al., 2021b) selects ${\mathbf{S}}_{j}$ of size $k$ via a mixed discrete–continuous bi-level objective:
@@ E.2.3 GraphCut-based Data Subset Selection (GraphCut)
GraphCut (Iyer et al., 2021) selects ${\mathbf{S}}_{j}$ via the generalized graph-cut function
@@ E.2.4 Reconstructing the Decision Boundary (BoundarySet-CCS)
BoundarySet-CCS (Yang et al., 2024) selects samples near the model’s decision boundary and then enforces coverage across distance bands. Distance-to-boundary is approximated per sample by the minimum number of PGD steps required to flip its prediction; CCS (coverage-centric sampling) then allocates the coreset budget across bands to preserve distribution coverage.

### v2

@@ 7 Computational and Storage Efficiency
Beyond accuracy, a practical coreset method should keep both compute and storage costs low. CLD does so by relying only on per-sample loss scalars that standard training already produces; no per-sample gradients, Hessians, or pairwise similarities are required. Practically, this entails scoring the full training set at a small number of checkpoints, but the scores are exactly the per-sample losses computed during training, i.e., no extra inference sweeps. The only extra work is a forward-only sweep over a small held-out query set each proxy epoch ($Q\ll N$) to record query losses. In contrast, gradient/adversarial methods incur extra backward passes, while similarity/nearest-neighbor methods require feature extraction and large feature caches. We quantify the end-to-end compute cost (selection plus training on the selected coreset) and the storage overhead in detail in Appendix˜E and summarize the symbolic complexity below.
@@ 8 Discussion
Several methods incorporate bias-reduction mechanisms, such as $\mathtt{CCS}$ (zheng_ccs), which stratifies selection across score percentiles to promote diversity. In contrast, $\mathtt{CLD}$ leverages per-class validation trajectories, yielding a generalization-aware signal that naturally balances classes and downweights noisy or redundant samples. As shown in Figure˜5(b), applying $\mathtt{CCS}$-style stratified sampling on top of $\mathtt{CLD}$ scores consistently reduces accuracy across coreset sizes on CIFAR-100. This contrasts with metrics like $\mathtt{AUM}$, where $\mathtt{CCS}$ can be beneficial; for $\mathtt{CLD}$, percentile quotas perturb its validation-aligned ranking and reintroduce less informative points.
@@
Our theoretical guarantees for $\mathtt{CLD}$ require that the validation set be a faithful proxy for the test distribution (Assumption˜3 in Section˜5). To understand this, we ask: How sensitive is $\mathtt{CLD}$ to the composition of the validation set, and does bias in this validation set affect downstream coreset quality? To probe this, we split CIFAR-100’s $50\text{k}$ training set into a classwise $25\text{k}/25\text{k}$ train/pool partition. From the pool, we constructed $5000$ example validation sets using five heuristics based on publicly available memorization scores (FZ_infl_feldman2020), denoted $\mathtt{mem}$. Intuitively, low-$\mathtt{mem}$ points correspond to canonical, stereotypical examples (often helpful for transfer and exploited by methods such as SloCurv), while high-$\mathtt{mem}$ points surface atypical or mislabeled examples. These heuristics, therefore, let us bias the validation set toward “typical” or “atypical” regions of the data. We trained ResNet-18 proxies on the $25\text{k}$ train split, computed $\mathtt{CLD}$ with respect to each heuristic-based validation set, built class-balanced coresets, and fine-tuned ResNet-18 across coreset sizes (five seeds; see Figure˜6(b)). As shown in Figure˜6(a), the non-random heuristics yield validation sets with markedly different $\mathtt{mem}$ profiles. Two clear patterns emerge. First, validation sets biased toward Highest-$\mathtt{mem}$ examples consistently degrade performance across coreset sizes: overrepresenting atypical or mislabeled samples lowers accuracy and hinders learning. Second, validation sets built from Lowest-$\mathtt{mem}$ examples generally support stronger generalization, but retaining some high-$\mathtt{mem}$ points is beneficial for capturing long-tail behavior likely present in the test distribution. In our runs, Proportional sampling, drawing examples according to the pool’s original $\mathtt{mem}$ distribution, was the most reliable overall, and would likely improve further if mislabeled points were filtered or downweighted (which we did not do in this study). These findings underscore that $\mathtt{CLD}$’s effectiveness depends critically on the quality of the validation signal: coresets cannot exceed the fidelity of the validation dynamics they are aligned with. This motivates future work on principled procedures to build clean, representative validation sets (e.g., robust to label noise) and to select validation sizes that balance reliability with efficiency.
@@
Loss Differences as a Gradient-Free Proxy for Influence.
The $\mathtt{CLD}$ metric is built on correlating per-example loss differences, a deliberate choice over simpler signals like raw losses or gradient norms. This choice is theoretically motivated: as we formalize in Lemma˜1, a first-order expansion shows that the loss difference $\Delta\ell(\theta^{t};z)$ approximates the gradient inner product $\langle\nabla_{\theta}\ell(\theta^{t-1};z),\,\delta\theta^{t-1}\rangle$. This insight positions $\mathtt{CLD}$ as an efficient proxy for the alignment dynamics tracked by influence methods that compute gradient similarities, such as TracIn (tracin_pruthi2020). The primary advantage of this proxy approach is computational efficiency. Methods like TracIn are powerful but expensive, requiring the storage and processing of high-dimensional per-sample gradients. $\mathtt{CLD}$ captures the same underlying alignment dynamics while avoiding this overhead entirely. Simpler signals, such as raw losses or gradient norms, are even cheaper but conceptually flawed, as they are confounded by signal drift or discard crucial directional information. Empirically, $\mathtt{CLD}$’s performance is comparable to TracIn on key attribution metrics, including the Linear Datamodeling Score (a measure of how well scores correlate with true sample influence) and prediction brittleness (a measure of how critical top-ranked samples are to model predictions). See Appendix˜F for the full analysis and a broader discussion of how $\mathtt{CLD}$ relates to influence methods.
Beyond Supervised Vision: Scope and Caveats.
Because $\mathtt{CLD}$ requires only per-example training losses and a small validation proxy, its core criterion can be adapted to other supervised settings (e.g., contrastive learning with a validation loss, or object detection by aggregating per-instance losses to a per-image score). In this work, we intentionally limit our scope to supervised image classification for controlled comparisons and do not claim empirical validation outside this domain. Applying $\mathtt{CLD}$ to modern language model fine-tuning, however, is non-trivial due to challenges like the “squeezing effect” (ren2025_learning). As fine-tuning progresses, the model concentrates probability mass onto the exact training sequences, which can paradoxically lower the assigned probabilities of similar but non-identical validation samples. This causes a divergence between training and validation loss trajectories, meaning that training samples promoting generalization are not guaranteed to have a high $\mathtt{CLD}$ score. Corroborating this finding, xia2024_LESS observe that, unlike in vision, minimizing validation loss does not reliably improve model performance in instruction tuning. Adapting a $\mathtt{CLD}$-style selector for LLMs will therefore require developing task-appropriate generalization proxies, which is an important direction for future work.
@@ Acknowledgements
This work was supported in part by the Center for the Co-Design of Cognitive Systems (CoCoSys), a DARPA-sponsored JUMP 2.0 center, the Semiconductor Research Corporation (SRC), the National Science Foundation, and Collins Aerospace. We are also thankful to Efstathia Soufleri, Akshita Gupta, Utkarsh Saxena, Amitangshu Mukherjee, and Sakshi Choudhary for their helpful discussions and feedback.
@@ 16  ${\mathbf{V}}_{c}\leftarrow\{\vec{q}_{j}\in{\mathbf{V}}:y_{j}=c\}$;
Implementation note. Loss values in Steps 3–6 are logged as part of the normal training loop; no additional passes are introduced.
@@ Appendix B Detailed Theoretical Framework
B.1 Roadmap and Notation
Proof Outline. Our main convergence guarantee, Theorem 1, is built upon three supporting lemmas:
  • Lemma 1, which establishes that a high $\mathtt{CLD}$ score implies strong alignment between a sample’s gradient and the validation gradient.
  • Lemma 2, which shows that this alignment is preserved during coreset training, provided the coreset and full-data parameter trajectories remain close.
  • Lemma 3, which leverages this stable alignment to bound the approximation error between the average coreset gradient and the true population risk gradient.
Notation Summary. We use the following symbols throughout our analysis:
  • $L$: The smoothness constant of the loss function.
  • $B$: An upper bound on the per-sample gradient norm.
  • $k$: The size of the coreset.
  • $Q$: The size of the validation set.
  • $\epsilon$: A parameter controlling the strictness of CLD-based sample selection.
  • $\kappa$: The alignment-gap term, quantifying the gradient mismatch from coreset selection.
  • $\delta$: The validation representativeness error, which decays as $O(\sqrt{Q})$.
  • $\eta$: The learning rate (step size) of the optimizer.
  • $T$: The total number of training iterations.
  • $\theta_{S}^{t},\theta_{C}^{t}$: Model parameters at iteration $t$ when training on the full dataset and the coreset, respectively.
B.2 Supporting Lemmas for Theorem˜1
@@
Proof Outline.
Use first-order loss changes to relate per-example loss differences and validation loss differences at consecutive steps. High correlation of differences implies an (approximately) positive linear link between their inner products with update directions, forcing the sample gradient to be a positive scalar multiple of the validation gradient under sufficiently varied updates. Continuity then yields $\cos\!\left(\angle(\nabla\ell(\cdot,\vec{z}_{m}),G_{{\mathbf{V}}})\right)\geq 1-\epsilon^{\prime}_{t}$ with $\epsilon^{\prime}_{t}\!\to\!0$ as $\epsilon\!\to\!0$. ∎
@@
Proof outline..
Compare gradients at $\theta_{{\mathbf{C}}}^{t}$ and $\theta_{{\mathbf{S}}}^{t}$ using $L$-smoothness to bound deviations by $O(\|\delta_{t}\|)$. Expand the inner product and bound cross-terms via Cauchy–Schwarz and the gradient-norm bound $B$. Plug the base alignment from Lemma 1 at $\theta_{{\mathbf{S}}}^{t}$ to transfer alignment to $\theta_{{\mathbf{C}}}^{t}$, yielding the stated $1-\kappa$ lower bound with linear/quadratic dependence on $\|\delta_{t}\|$. ∎
@@
Proof outline..
Split the error as $\|\gamma_{{\mathbf{C}}}^{t}-G_{{\mathbf{V}}}\|+\|G_{{\mathbf{V}}}-\nabla R_{{\mathbf{D}}}\|$. The second term is $\delta$ by validation representativeness. For the first, average the per-sample deviations and apply Jensen to get a mean of squared distances; combine with Lemma 2 to bound each by $2B^{2}\kappa$, yielding $\|\gamma_{{\mathbf{C}}}^{t}-G_{{\mathbf{V}}}\|\leq B\sqrt{2\kappa}$. ∎
@@
B.3 Proof for Theorem˜1
Proof outline..
Apply the $L$-smooth descent lemma to $R_{\mathbf{D}}(\theta)$ with the update $\theta_{{\mathbf{C}}}^{t+1}=\theta_{{\mathbf{C}}}^{t}-\eta\,\gamma_{{\mathbf{C}}}^{t}$ to obtain a one-step inequality involving $\langle\nabla R_{{\mathbf{D}}},\gamma_{{\mathbf{C}}}^{t}\rangle$. Decompose $\gamma_{{\mathbf{C}}}^{t}=\nabla R_{{\mathbf{D}}}+(\gamma_{{\mathbf{C}}}^{t}-\nabla R_{{\mathbf{D}}})$ and bound the error term using Lemma 3: $E_{t}\!=\!\|\gamma_{{\mathbf{C}}}^{t}-\nabla R_{{\mathbf{D}}}\|\!\leq\!B\sqrt{2\kappa}+\delta$. Control the mixed term by Cauchy–Schwarz + Young; bound $\|\gamma_{{\mathbf{C}}}^{t}\|$ by $B$. Sum over $t$ to telescope $R_{t}-R_{t+1}$ and isolate $\min_{t}\|\nabla R_{{\mathbf{D}}}(\theta_{{\mathbf{C}}}^{t})\|^{2}$, yielding the stated bound with $(B\sqrt{2\kappa}+\delta)^{2}$ and an $L\eta B^{2}$ residual. ∎
@@ Appendix C Datasets, Models, and Experimental Details
We evaluate our method on two standard image classification benchmarks: CIFAR-100 (cifar) and ImageNet-1k (imagenet). CIFAR-100 consists of 50,000 training and 10,000 test images across 100 classes and is publicly available without licensing restrictions. For ImageNet-1k, we use the official release from https://image-net.org/download.php, which is provided under a standard academic research license and requires user agreement to the terms of access.
@@
For baseline comparisons, we use the DeepCore library deepcore (https://github.com/PatrickZH/DeepCore), which provides standardized implementations of several coreset selection techniques and is licensed under MIT. All training and evaluation code was implemented in PyTorch; dependencies including torchvision are MIT/BSD licensed.
@@
All networks were trained using SGD sgd for $164$ epochs with an initial learning rate of $0.1$, decayed by a factor of $0.1$ at epochs $81$ and $121$. Nesterov momentum nesterov with momentum $0.9$ was used, along with weight decay $5\times 10^{-4}$. Standard augmentations included resizing to $32\times 32$, random cropping with padding =4=4, random horizontal flips, and normalization.
@@ Appendix E Detailed Explanation of Compute and Storage Cost of Coreset Methodologies
  • We use the standard ImageNet recipe of $T{=}T_{\text{proxy}}{=}90$ epochs (bearpaw_github).
@@ E.1.1 Kernel Herding (Herding)
Herding (herding_chen2010) iteratively constructs a representative subset by approximating the data distribution in an RKHS. At iteration $t$, it selects
@@ E.1.2 Example Forgetting (Forgetting)
Forgetting (forgetting_toneva2018) measures, for each training sample, how many times it transitions from being correctly classified to incorrectly classified during training (“forgetting events”). Examples with higher forgetting counts are ranked as more informative.
@@ E.1.3 Area Under Margin (AUM)
AUM (pleiss2020_aum) scores each training sample by aggregating its margin over training (e.g., logit of the true class minus the max non-true logit), producing the area under the margin across epochs/updates. Higher absolute AUM indicates more consistently confident predictions; lower AUM can flag ambiguous or noisy samples. We compute AUM with a proxy model and then train the large model on the selected coreset.
@@ E.1.4 Contrastive Active Learning (Cal)
Cal (cal_margatina2021) acquires unlabeled examples that are near labeled ones in feature space yet differ in predictive probabilities (contrastive pairs), using nearest neighbors over encoder features and a simple divergence-based ranking.
@@ E.1.5 Gradient and Error L2 Norm-based Data Pruning (GraNd, EL2N)
GraNd (E2LN_grand_paul2021) ranks training examples by the (expected) per-example gradient norm early in training:
@@
EL2N (E2LN_grand_paul2021) ranks examples by the (expected) L2 error of predictions early in training:
@@ E.1.6 Using Class Feature Medians (Moderate)
Moderate (xia2022moderate) builds a representative coreset by selecting, within each class, the samples whose feature-to-center distances are closest to that class’s median distance (thus avoiding both easy near-center redundancies and far-out outliers). We compute class centers and distances in a proxy feature space.
@@ E.1.7 Message Passing ($\mathbb{D}^{2}-\mathtt{Pruning}$)
$\mathbb{D}^{2}\mathtt{Pruning}$ (modelpred_d2_maharana2023) selects a coreset by balancing difficulty and diversity via message passing on a dataset graph built from proxy features. Initial per-sample difficulty scores (from the proxy) are diffused over a kkNN graph so that each example’s score incorporates information from its neighbors; a graph-based sampler then selects a subset that covers diverse yet difficult regions.
@@ E.2.1 Gradient Matching Optimization (CRAIG)
CRAIG (craig_mirzasoleiman2020) selects a subset whose (aggregated) gradients closely match those of the full dataset across training, typically using a submodular (stochastic-greedy) objective over per-sample gradient embeddings at selected anchor epochs. We compute embeddings with a proxy model and then train the large model on the coreset for all $T$ epochs.
@@ E.2.2 Generalization-based Data Subset Selection for Efficient and Robust Learning (Glister)
Glister (glister_killamsetty2021) selects ${\mathbf{S}}_{j}$ of size $k$ via a mixed discrete–continuous bi-level objective:
@@ E.2.3 GraphCut-based Data Subset Selection (GraphCut)
GraphCut (graphcut_iyer2021) selects ${\mathbf{S}}_{j}$ via the generalized graph-cut function
@@ E.2.4 Reconstructing the Decision Boundary (BoundarySet-CCS)
BoundarySet-CCS (mindboundary_yang2024) selects samples near the model’s decision boundary and then enforces coverage across distance bands. Distance-to-boundary is approximated per sample by the minimum number of PGD steps required to flip its prediction; CCS (coverage-centric sampling) then allocates the coreset budget across bands to preserve distribution coverage.